Shift report 20011019 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Fri Oct 19 2001 - 08:05:11 EDT

  • Next message: HAGEL@comp.tamu.edu: "Last chance to comment on "final" version of NIM paper"

    ------------------------------------------------------------
    Date: Fri Oct 19 08:05:11 2001, Shift: 00:00-08:00
    Supervisor: KH, with: 
    Log book#10 pages 32 - 34
    
    Shift Summary:
    ------------------------------------------------------------
    A good store from midnight to 4:45 when BRAHMS + PHENIX
    voted to end store.
    A new store started in good time, but the ramp failed.
    Shortly after that, the monitor said:
    "Power Supply Control problem.  An expert is investigating."
    At the end of the shift, there is a Cryogenic System problem.
    
    Runs achieved:
    run     trig1     trig 6
    5435     401662   128359
    5436     -----     30000 estimated
    5438     122137    38803
    
    Run 5436 ran for 45 minutes, then crashed with an IoServer
    error.  It took one false start, but I got it going again
    by stopping IoServer, restarting RunServer and then
    starting IoServer again.  Because of that, the trigger
    entries were not logged into the rundb.
    
    The end of run 5438 brings the total trig 6 triggers to
    387958.  After the store was over, I therefore had MCR
    come and change the polarities.  So the polarities of
    the FS magnets are now set on B and the polarity of D5 is
    now A.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Fri Oct 19 2001 - 14:06:00 EDT